# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ Abstract
We derive the beam tracing and profile evolution for the propagation of any localised beam with arbitrary profile through an inhomogeneous cold plasma. We recover standard Gaussian beam-tracing, with an additional PDE describing the evolution of the beam’s profile as it propagates through the plasma. We then solve for generic families of solutions to the PDE using ladder operators, which can be chosen to reduce to Gauss-Hermite beams in homogeneous media. We importantly obtain an exact expression for the resulting beam profile, demonstrating that Hermite modes will generally evolve into a superposition of different modes during propagation through inhomogeneous plasmas, contrary to prior work on the subject. Importantly, this approach allows us to construct an inner product with orthogonality between solutions for the beam evolution, a useful feature for future analysis of the diagnostic signal received from arbitrary beams.
@@ I Introduction
The propagation of electromagnetic beams through plasmas has long been an important aspect of nuclear fusion research. The ability to model and interpret the evolution of a beam’s profile and power deposition has proven to be important parts of our plasma control toolbox, via heating, current driveErckmann1994; wolf2018; sysoeva2015, and even suppression of MHD instabilitiesLaHaye2009. On top of this, microwave diagnostic techniques are versatile tools for real-time plasma diagnostics. With appropriate engineering, they can be deployed in environments with high neutron fluxGoncalves2023engineering, as well as provide a broad range of measurements – we diagnose electron temperature via electron-cyclotron emission(ECE)bornatici1983ece; classen2010ece; liu2018ece, electron density via Doppler reflectometryHillesheim2009; Conway2004; gusakov2004; Lin2024reflectometry, and line-integrated magnetic information via polarimetrysmith2008polarimetry; zhu2015polarimetry; chen2018polarimetry. Notably, techniques like Doppler backscattering(DBS)HallChen2022; HallChen2022b; HallChen2024; liang2025; yeoh2025; RuizRuiz2025; Hennequin2006 and cross-polarisation scattering(CPS)stepanov1998; colas1996 further enable studies of electromagnetic turbulence and MHD activity with higher temporal and spatial resolution. On top of this, many of the aforementioned systems are capable of measuring even more plasma properties with modern advancementsKohn2025review. Given the incredible robustness of these microwave techniques, accurate modelling of microwave interactions and measurements is essential.
@@
Historically, ray-tracing codessmirnov2001genray; Esterkin1996; xie2022boray are typically sufficient for understanding the propagation of a beam through a plasma, by performing ray tracing across multiple(possibly coupled) rays to model how the beam profile might evolve. Beam tracing offers faster and more accurate simulations by modelling the evolution of the profile of the beam itself. This allows us to model diffraction effects with greater accuracy compared to a simplistic ray description of a beam. TORBEAMPoli2001; Poli2018, which was built upon theoretical work by PereversevPere1998, was one of the first codes developed for modelling beam evolution using beam tracing. Although it certainly set the groundwork for synthetic diagnostic and heating techniques, newer codes like ScottyHallChen2022; Wikarta2025, WKBeamweber2015; Snicker2018 and PARADEPARADE2021 have since been developed and further optimised for their specific use cases, instead of being one-size-fits-all codes like TORBEAM.
@@
Given the prevalence and pertinence of synthetic beam tracing applications in plasma diagnostics and control, we revisit the propagation of an arbitrary beam profile in an anisotropic plasma. Whilst the propagation of the lowest order Gauss-Hermite mode(a Gaussian beam) has been well understood, the higher order modes are still not that well understood. We take a different approach to previous work, by directly solving for the evolution of an arbitrary beam profile. In this manuscript, we focus on solving for the propagation of electromagnetic beams in cold plasmas exactly(within certain orderings), which will ultimately be most applicable to the regime of microwave diagnostic techniques. Our goal is not only to model the propagation of microwave beams in general, but to set the theoretical groundwork for improved accuracy in signal analysis, by modelling the evolution of non-Gaussian imperfections in the antenna pattern.
@@
Previous work by PereversevPere1998 obtained a set of beam-tracing equations that specifically modeled the evolution of Gauss-Hermite modes, which in principle allows one to model the evolution of any beam. However, that was done by directly substituting a Gauss-Hermite ansatz into the Helmholtz equation. Pereversev finds a physicist’s Hermite polynomial basis of solutions to the beam evolution, which he crucially argues is possible through eqs. 3.12-3.13. What we discover is that in general, beams that start out as Gauss-Hermite beams may evolve into some other kind of profile, which we ascertained exactly to be a linear combination of Hermite modes in a rotated basis [^1] It matches with the orthonormal basis generated from Pereversev’s basis using Gram-Schmidt decomposition on his coordinate system, and the vectors we use for said decomposition match exactly with his basis., contrary to his results. We discuss this discrepancy in greater detail in section V.2. Our work has the advantage of providing immense freedom to generate many different bases of solutions, along with being able to model the evolution of each basis of solutions exactly. This should allow one to choose bases most suited for their problem at hand.
@@
We first derive beam-tracing for any arbitrary beam using perturbation theory up to the second order in an inhomogeneous cold plasma. We make a distinction between the Gaussian beam envelope and the beam profile, which is some function attenuated by said envelope. Expectedly, we show that the exact distinction is not a matter for concern, as the exact choices for each do not affect the consistency of our results. We discover that the evolution of our beam is governed by the same set of beam-tracing equations for a Gaussian beam envelope as in prior, separate work by Pereversev and Hall-Chen et al.HallChen2022, with an added second-order partial differential equation governing the evolution of the beam profile. The profile evolution equation we obtain demonstrates that Hermite polynomial beam profiles are generally not preserved in an inhomogeneous plasma, although a constant profile expectedly remains unchanged.
@@
Having to solve a PDE ultimately defeats the purpose of this approach, as the main advantage of the beam-tracing approach is to avoid the need to solve any PDEs. We thus proceed to solve the PDE in general, using ladder operators to generate two infinite, biorthogonal families of solutions. In particular, we focus on constructing ladder operators that reduce to Gauss-Hermite polynomials at the boundary of the plasma. Not only that, we confirm that Gauss-Hermite polynomials are solutions in homogeneous, isotropic media, expectedly in agreement with the conventional wisdom. In general, the solutions do not have to be Hermite polynomials at the plasma boundary, and there are many other solutions that could be utilised. However, choosing a biorthogonal set of Hermite solutions at the boundary actually gives us a biorthogonal set of solutions throughout, a result we will prove using our ladder operators. Crucially, this is used to define an orthogonal inner-product on the Hermite-like basis of solutions. This result is also generally applicable in the broader context of mathematical modelling, particularly in the theory of the time-dependent Ornstein-Uhlenbeck processVatiwutipong2019. Finally, we incorporate weak dissipative effects for modelling microwave heating of the plasma.
@@
We finally discuss potential future applications of our work in microwave diagnostics.
@@ II Beam Tracing Equations
In this section, we lay out the derivation of our beam-tracing equations, utilising perturbation theory. To perform the expansion, we will order that $\frac{\lambda}{W}\sim\frac{W}{L}<<1$. $\lambda$ refers to the wavelength of our beam, W refers to the width of our beam, and L refers to the typical length scale of inhomogeneity in our plasma.
Our goal is to find the beam solution to the Helmholtz equation in a plasma. Repeated indices will be summed over whenever present in this manuscript. Setting $\frac{c}{\Omega}=1$, the Helmholtz equation is:
$(\delta_{il}\delta_{jm}-\delta_{im}\delta_{jl})\partial_{j}\partial_{l}(E_{m})=\epsilon_{ij}E_{j}$ (1)
We generally will not need to reintroduce $\frac{c}{\Omega}$ later on, as it turns out that the units of all later equations are already correct without it. $E_{i}$ refers to the components of the electric field at any point, and $\epsilon_{ij}$ is the cold plasma dielectric tensor. Note that we will be utilising the Euclidean metric for all summations, thus there is no difference between upper and lower indices, and they will be used as needed to keep the notation readable.
@@ II.1 Beam Ansatz
The beam ansatz we define only requires our beam to be localised in space, but it can have any cross section within that localised envelope. We need to define beam-centred coordinates. Suppose that the coordinates of our central ray are given by $\bm{q}(\tau)$ where $\tau$ simply parametrises the path. Then the ray velocity at every point is given by $\bm{g}=g\hat{\bm{g}}=\frac{\mathrm{d}\bm{q}}{\mathrm{d}\tau}$. $\bm{w}$ is the width perpendicular to said ray velocity vector at any point.
We thus adopt the following beam ansatz:
$\bm{E}(\bm{r})=\bm{A}\left(\tau,w_{x},w_{y}\right)\exp\left(\mathrm{i}\left(s(\tau)+\bm{K}_{w}\cdot\bm{w}+\frac{1}{2}\bm{w}\cdot\bm{\Psi}_{w}\cdot\bm{w}\right)\right),$ (2)
where $w_{x}$ and $w_{y}$ are components of $\bm{w}$(for the sake of localisation). $\bm{\Psi}_{w}$ is a complex matrix, where the imaginary part encodes the beam envelope while the real part encodes the curvature of the beam. $s(\tau)=\int_{0}^{\tau}K_{g}(\tau^{\prime})g\mathrm{d}\tau^{\prime}$, such that together with $\bm{K_{w}}$ we have a well defined wavevector $\bm{K}$ for our beam. Although we could always absorb the envelope and phase into the amplitude function $\bm{A}$, keeping it separate yields simpler and more illuminating mathematics. These quantities we have defined have the following orderings: $\bm{K}\sim\frac{1}{\lambda}$ and $\bm{\Psi}_{w}\sim\frac{1}{W^{2}}$. Thus, $s\sim\frac{L}{\lambda}$ and $\bm{K}_{w}\cdot\bm{w}\sim\frac{W}{\lambda}$.
@@
$\bm{A}\left(\tau,w_{x},w_{y}\right)=\mathcal{P}\left(\tau,w_{x},w_{y}\right)\alpha^{(0)}(\tau)\hat{\bm{e}}^{(0)}(\tau)+\bm{A}^{(1)}\left(\tau,w_{x},w_{y}\right)+\bm{A}^{(2)}\left(\tau,w_{x},w_{y}\right)+$... (3)
The superscripts indicate the order in $\frac{\lambda}{W}$ of the associated quantity. To lowest order, we have chosen to decompose our amplitude into a profile function $\mathcal{P}$ as well as an amplitude $\alpha^{(0)}$. This choice is primarily for compatibility with previous results and simplicity, but it also makes classifying beams with different profiles easier later on. It is important that our chosen profile must have weaker growth than our Gaussian beam envelope at all times, in order to be a localised beam.
@@
As demonstrated in previous workHallChen2022, these quantities are given by:
@@
where $\bm{\kappa}$ is the ray curvature, given by $\bm{\kappa}=\frac{1}{g}\frac{\mathrm{d}\hat{\bm{g}}}{\mathrm{d}\tau}\sim\frac{1}{L}$. In general, all derivatives along the ray length would be on the length scale of inhomogeneity of our plasma.
We now substitute our ansatz into the Helmholtz equation. We will use $\bm{E}=\bm{A}\exp(\mathrm{i}\phi)$. The right hand side of the Helmholtz equation can be expanded about the center of the ray, at any point along the ray:
$\epsilon_{ij}(\bm{r})=\epsilon_{ij}(\bm{q}(\tau))+w_{k}\partial_{k}\epsilon_{ij}(\bm{q}(\tau))+\frac{1}{2}w_{k}w_{l}\partial_{k}\partial_{l}\epsilon_{ij}(\bm{q}(\tau))+$... (6)
@@
We thus require all these various derivatives, along with their expansions and orderings. For the derivatives of the phase, using the fact that:
$\frac{\mathrm{d}\bm{K}_{w}}{\mathrm{d}\tau}=\left(\frac{\mathrm{d}K_{x}}{\mathrm{d}\tau}+K_{y}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)\hat{\bm{x}}+\left(\frac{\mathrm{d}K_{y}}{\mathrm{d}\tau}+K_{x}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\right)\hat{\bm{y}}+\mathrm{terms\quad along\quad g},$ (8)
@@
$\begin{split}\frac{\mathrm{d}\bm{\Psi}_{w}}{\mathrm{d}\tau}&=\left(\frac{\mathrm{d}\Psi_{xx}}{\mathrm{d}\tau}+2\Psi_{xy}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)\hat{\bm{x}}\hat{\bm{x}}\\ &+\left(\frac{\mathrm{d}\Psi_{xy}}{\mathrm{d}\tau}+\Psi_{xx}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}+\Psi_{yy}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)(\hat{\bm{x}}\hat{\bm{y}}+\hat{\bm{y}}\hat{\bm{x}})\\ &+\left(\frac{\mathrm{d}\Psi_{yy}}{\mathrm{d}\tau}+2\Psi_{xy}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\right)\hat{\bm{y}}\hat{\bm{y}}+\mathrm{terms\quad along\quad g},\end{split}$ (9)
@@
$\bm{\nabla}\phi=\bm{K}+\frac{\hat{\bm{g}}}{g(1-\bm{\kappa}\cdot\bm{w})}\left(\frac{\mathrm{d}\bm{K}}{\mathrm{d}\tau}\cdot\bm{w}\right)+\bm{\Psi}_{w}\cdot\bm{w}+\frac{\hat{\bm{g}}}{g(1-\bm{\kappa}\cdot\bm{w})}\left(\frac{1}{2}\bm{w}\cdot\frac{\mathrm{d}\bm{\Psi}_{w}}{\mathrm{d}\tau}\cdot\bm{w}\right).$ (10)
Thus, to get the various orders of $\bm{\nabla}\phi$, we simply have to expand $\frac{1}{(1-\bm{\kappa}\cdot\bm{w})}=1+\bm{\kappa}\cdot\bm{w}+(\bm{\kappa}\cdot\bm{w})^{2}+$....
As for the second order derivatives, we will only require $(\partial_{i}\partial_{j}\phi)^{(0)}$. We summarise all of the derivatives of the phase required in our derivation below:
@@
We have defined a generalised form of $\bm{\Psi}_{w}$ called $\bm{\Psi}$, as the projection onto the width of said quantity will provide us with precisely $\bm{\Psi}_{w}$. $\bm{\Psi}$ will make our subsequent mathematics much simpler.
Next, we wish to obtain the ordering for the derivatives of our amplitude function. We once again summarise any necessary derivatives below:
@@
One can notice that $\bm{\nabla}\phi^{(n)}\sim\frac{1}{\lambda}(\frac{\lambda}{W})^{n}$ whilst $(\bm{\nabla}A_{m})^{(n)}\sim\frac{\mathrm{Electric\,field}}{W}(\frac{\lambda}{W})^{n}$. The derivatives of various orders of our profile function are given by:
$\begin{split}\bm{\nabla}(\mathcal{P})^{(0)}&=\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P},\\ \bm{\nabla}(\mathcal{P})^{(1)}&=\left(\frac{\hat{\bm{g}}}{g}\frac{\partial}{\partial\tau}+\frac{w_{y}\hat{\bm{g}}}{g}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\frac{\partial}{\partial w_{x}}+\frac{w_{x}\hat{\bm{g}}}{g}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P},\\ \bm{\nabla}\bm{\nabla}(\mathcal{P})^{(0)}&=\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P}.\end{split}$ (13)
Note that our ordering here requires:
$\frac{|\bm{\nabla}\mathcal{P}^{(0)}|}{\mathcal{P}}\sim\frac{1}{W}\quad\mathrm{and}\quad\frac{|\frac{\partial\mathcal{P}}{\partial\tau}|}{\mathcal{P}}\sim\frac{g}{L}.$ (14)

### v3

@@ Abstract
We derive the beam tracing and profile evolution for the propagation of any localised beam with arbitrary profile through an inhomogeneous cold plasma. We recover standard Gaussian beam-tracing, with an additional PDE describing the evolution of the beam’s profile as it propagates through the plasma. We then solve for generic families of solutions to the PDE using ladder operators, which can be chosen to reduce to Gauss-Hermite beams in homogeneous media. We importantly obtain an exact expression for the resulting beam profile, demonstrating that Hermite modes will generally evolve into a superposition of different modes during propagation through inhomogeneous plasmas, contrary to prior work on the subject. Importantly, this theory provides us with a lightweight numerical approach for modelling microwave beam propagation through a plasma, which is easily implemented on top of existing Gaussian beam tracing techniques.
@@ I Introduction
The propagation of electromagnetic beams through plasmas has long been an important aspect of nuclear fusion research. The ability to model and interpret the evolution of a beam’s profile and power deposition has proven to be important parts of our plasma control toolbox, via heating, current drive[Erckmann1994, wolf2018, sysoeva2015], and even suppression of MHD instabilities[LaHaye2009]. On top of this, microwave diagnostic techniques are versatile tools for real-time plasma diagnostics. With appropriate engineering, they can be deployed in environments with high neutron flux[Goncalves2023engineering], as well as provide a broad range of measurements – we diagnose electron temperature via electron-cyclotron emission(ECE)[bornatici1983ece, classen2010ece, liu2018ece], electron density via Doppler reflectometry[Hillesheim2009, Conway2004, gusakov2004, Lin2024reflectometry], and line-integrated magnetic information via polarimetry[smith2008polarimetry, zhu2015polarimetry, chen2018polarimetry]. Notably, techniques like Doppler backscattering(DBS)[HallChen2022, HallChen2022b, HallChen2024, liang2025, yeoh2025, RuizRuiz2025, Hennequin2006] and cross-polarisation scattering(CPS)[stepanov1998, colas1996] further enable studies of electromagnetic turbulence and MHD activity with higher temporal and spatial resolution. On top of this, many of the aforementioned systems are capable of measuring even more plasma properties with modern advancements[Kohn2025review]. Given the incredible robustness of these microwave techniques, accurate modelling of microwave interactions and measurements is essential.
@@
Historically, ray-tracing codes[smirnov2001genray, Esterkin1996, xie2022boray] are typically sufficient for understanding the propagation of a beam through a plasma, by performing ray tracing across multiple(possibly coupled) rays to model how the beam profile might evolve. Beam tracing offers faster and more accurate simulations by modelling the evolution of the profile of the beam itself. This allows us to model diffraction effects with greater accuracy compared to a simplistic ray description of a beam. TORBEAM[Poli2001, Poli2018], which was built upon theoretical work by Pereversev[Pere1998], was one of the first codes developed for modelling beam evolution using beam tracing. It has certainly set the groundwork for synthetic diagnostic and heating techniques, while newer codes like Scotty[HallChen2022, Wikarta2025], WKBeam[weber2015, Snicker2018] and PARADE[PARADE2021] have since been developed and further optimised for their specific use cases, instead of being one-size-fits-all codes like TORBEAM.
@@
Given the prevalence and pertinence of synthetic beam tracing applications in plasma diagnostics and control, we revisit the propagation of an arbitrary beam profile in an anisotropic plasma. Whilst the propagation of the lowest order Gauss-Hermite mode(a Gaussian beam) has been well understood, the higher order modes are still not that well understood. Pereversev’s early theoretical work on the subject utilised asymptotic expansion of a small parameter $\sqrt{\kappa}$, similar to the parameter $\frac{\lambda}{W}$ we will be using in this work. The issue with this approach is that these orderings break down near a caustic, making them unreliable near caustics. On the other hand, metaplectic geometrical optics(MGO) is the most modern approach currently in use. It generalizes beam tracing whilst remaining valid through caustics and mode-conversion regions[lopez2022metaplectic, lopez2024, donnelly2021steepest, hojlund2024metademo]. In the paraxial limit, MGO reduces to standard beam tracing. Despite this, we focus on solving for the propagation of electromagnetic beams in cold plasmas exactly using the asymptotic expansion approach. This approach is mathematically more tractable than MGO, and still provides us with important insights away from caustics. Our goal is not only to model the propagation of microwave beams in general, but to set the theoretical groundwork for improved accuracy in signal analysis, by modelling the evolution of non-Gaussian imperfections in the antenna pattern.
@@
Whilst it may seem that this would offer nothing new that Pereversev’s approach did not already show, we surprisingly discover that beams that start out as Gauss-Hermite beams may evolve into some other kind of profile, which we ascertained exactly to be a linear combination of Hermite modes in a rotated basis. This is in contrast to Pereversev’s results, suggesting that this approach has not yet been well-understood and deserves more attention. We discuss this potential discrepancy in greater detail in section VI. Apart from simple application to a Gauss-Hermite basis, our work has the advantage of providing immense freedom to generate many different bases of solutions, which allows us to model the evolution of each basis of solutions exactly. This should allow one to choose a basis most suited for their problem at hand, for example, when choosing between ‘elegant’ and ‘traditional’ Hermite modes in Gaussian wave optics[pampaloni2004].
@@
We first derive beam-tracing for any beam using asymptotic expansion up to the second order in an inhomogeneous cold plasma. We make a distinction between the Gaussian beam envelope and the beam profile, which is some function attenuated by said envelope. This is purely a mathematical choice. We demonstrate in the case of a Gaussian beam that the exact distinction of ‘envelope’ and ‘profile’ do not affect the consistency of our results, as one would expect. We discover that the evolution of our beam is governed by the same set of beam-tracing equations for a pure Gaussian beam envelope as in prior, separate work by Pereversev and Hall-Chen et al.[HallChen2022], with an added second-order partial differential equation governing the evolution of the beam profile. The profile evolution equation we obtain demonstrates that Hermite polynomial beam profiles are generally not preserved in an inhomogeneous plasmas, although a constant profile expectedly remains unchanged.
@@
We proceed to solve the PDE in general, using ladder operators to generate two infinite, biorthogonal families of solutions. In particular, we focus on constructing ladder operators that reduce to Gauss-Hermite polynomials at the plasma-vacuum boundary. We confirm that Gauss-Hermite polynomials are solutions in homogeneous, isotropic media, in agreement with the conventional wisdom. In general, the solutions do not have to be Hermite polynomials at the plasma boundary, and there are many other solutions that could be utilised. Finally, we incorporate weak dissipative effects for modelling microwave heating of the plasma.
@@
We finally discuss potential future implementation and applications of our work in microwave diagnostics.
@@ II Beam Tracing Equations
In this section, we lay out the derivation of our beam-tracing equations, utilising perturbation theory. To perform the expansion, we will order that ${\lambda}/{W}\sim{W}/{L}\ll 1$. $\lambda$ refers to the wavelength of our beam, W refers to the width of our beam, and L refers to the typical length scale of inhomogeneity in our plasma.
Our goal is to find the beam solution to the Helmholtz equation in a plasma. Repeated indices will be summed over whenever present in this manuscript. Occasionally, Einstein notation is utilised when it presents greater simplicity than tensor notation. We are ultimately interested in finding solutions to the Helmholtz equation for an electromagnetic wave, which is given by the following in a plasma:
$\frac{c^{2}}{\Omega^{2}}\bm{\nabla}\times(\bm{\nabla}\times\bm{E})=\bm{\epsilon}\cdot\bm{E}.$ (1)
$\bm{E}$ refers to the components of the electric field at any point without the time factor $\exp(-\mathrm{i}\omega t)$, and $\bm{\epsilon}$ is the cold plasma dielectric tensor. Note that we will be utilising the Euclidean metric for all summations, thus there is no difference between upper and lower indices, and they will be used as needed to keep the notation readable.
@@ II.1 Beam Ansatz
The beam ansatz we define only requires our beam to be localised in space, but it can have any cross section within that localised envelope. The coordinates of our central ray are given by $\bm{q}(\tau)$ where $\tau$ simply parametrises the path. The ray velocity at every point is given by $\bm{g}=g\hat{\bm{g}}={\mathrm{d}\bm{q}}/{\mathrm{d}\tau}$. $\bm{w}$ is the width perpendicular to said ray velocity vector at any point. We thus adopt the following beam ansatz:
$\begin{split}\bm{E}(\bm{r})&=\bm{A}\left(\tau,w_{x},w_{y}\right)\exp\left(\mathrm{i}\psi\right),\\ \psi&=s(\tau)+\bm{K}_{w}\cdot\bm{w}+\frac{1}{2}\bm{w}\cdot\bm{\Psi}_{w}\cdot\bm{w},\end{split}$ (2)
[…]
@@
$\bm{A}\left(\tau,w_{x},w_{y}\right)=\mathcal{P}\left(\tau,w_{x},w_{y}\right)A^{(0)}(\tau)\hat{\bm{e}}^{(0)}(\tau)+\bm{A}^{(1)}\left(\tau,w_{x},w_{y}\right)+\bm{A}^{(2)}\left(\tau,w_{x},w_{y}\right)+$... (3)
The superscripts indicate the order in $\frac{\lambda}{W}$ of the associated quantity. To lowest order, we have chosen to decompose our amplitude into a profile function $\mathcal{P}$ as well as an amplitude $A^{(0)}$. This choice is primarily for compatibility with previous results and simplicity, but it also makes classifying beams with different profiles easier later on. It is important that our chosen profile must have weaker growth than our Gaussian beam envelope at all times, in order to be a localised beam.
@@
As derived in previous work[HallChen2022], these quantities are given by:
@@
where $\bm{\kappa}$ is the ray curvature, given by $\bm{\kappa}=({1}/{g})\,({\mathrm{d}\hat{\bm{g}}}/{\mathrm{d}\tau})\sim{1}/{L}$. In general, all derivatives along the ray length would be on the length scale of inhomogeneity of our plasma. Using this our, gradient operator thus has the following form:
$\bm{\nabla}=\underbrace{\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}}_{\bm{\nabla}_{w}}+\frac{\hat{\bm{g}}}{g(1-\bm{\kappa}\cdot\bm{w})}\left(\frac{\partial}{\partial\tau}-\frac{\mathrm{d}\bm{w}}{\mathrm{d}\tau}\cdot\bm{\nabla}_{w}\right)$ (6)
We now substitute our ansatz in eq. 2 into the Helmholtz equation. We will use $\bm{E}=\bm{A}\exp(\mathrm{i}\psi)$. The right hand side of the Helmholtz equation can be expanded about the center of the ray, at any point along the ray:
$\bm{\epsilon}(\bm{r})=\bm{\epsilon}(\bm{q}(\tau))+\bm{w}\cdot\bm{\nabla}\bm{\epsilon}(\bm{q}(\tau))+\frac{1}{2}\bm{ww}:\bm{\nabla}\bm{\nabla}\bm{\epsilon}(\bm{q}(\tau))+$... (7)
@@
We thus require all these various derivatives, along with their expansions and orderings. For the derivatives of the phase $\psi$, using the fact that:
$\left(\frac{\mathrm{d}\bm{K}_{w}}{\mathrm{d}\tau}\right)_{w}=\left(\frac{\mathrm{d}K_{x}}{\mathrm{d}\tau}+K_{y}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)\hat{\bm{x}}+\left(\frac{\mathrm{d}K_{y}}{\mathrm{d}\tau}+K_{x}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\right)\hat{\bm{y}}$, (9)
@@
[…]
@@
$\bm{\nabla}\psi=\bm{K}+\frac{\hat{\bm{g}}}{g(1-\bm{\kappa}\cdot\bm{w})}\left(\frac{\mathrm{d}\bm{K}}{\mathrm{d}\tau}\cdot\bm{w}\right)+\bm{\Psi}_{w}\cdot\bm{w}+\frac{\hat{\bm{g}}}{g(1-\bm{\kappa}\cdot\bm{w})}\left(\frac{1}{2}\bm{w}\cdot\frac{\mathrm{d}\bm{\Psi}_{w}}{\mathrm{d}\tau}\cdot\bm{w}\right).$ (11)
Thus, to get the various orders of $\bm{\nabla}\psi$, we simply have to expand $(1-\bm{\kappa}\cdot\bm{w})^{-1}=1+\bm{\kappa}\cdot\bm{w}+(\bm{\kappa}\cdot\bm{w})^{2}+$..., as each term is an order ${\lambda}/{W}$ lower than the previous. As for the second order derivatives, we will only require $(\bm{\nabla}\bm{\nabla}\psi)^{(0)}$. We summarise all of the derivatives of the phase required in our derivation below:
@@
We have defined a generalised form of $\bm{\Psi}_{w}$ called $\bm{\Psi}$, as the projection onto the width of said quantity will provide us with precisely $\bm{\Psi}_{w}$. Next, we wish to obtain the ordering for the derivatives of our amplitude function. We once again summarise any necessary derivatives below:
@@
One can notice that $\bm{\nabla}\psi^{(n)}\sim\lambda^{-1}({\lambda}/{W})^{n}$ whilst $(\bm{\nabla}\bm{A})^{(n)}\sim({A^{(0)}}/{W})({\lambda}/{W})^{n}$. The derivatives of various orders of our profile function are given by:
$\begin{split}(\bm{\nabla}\mathcal{P})^{(0)}&=\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P}=\bm{\nabla}_{w}\mathcal{P},\\ (\bm{\nabla}\mathcal{P})^{(1)}&=\left(\frac{\hat{\bm{g}}}{g}\frac{\partial}{\partial\tau}+\frac{w_{y}\hat{\bm{g}}}{g}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\frac{\partial}{\partial w_{x}}+\frac{w_{x}\hat{\bm{g}}}{g}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P},\\ &=\frac{\hat{\bm{g}}}{g}\left(\frac{\partial}{\partial\tau}-\frac{\mathrm{d}\bm{w}}{\mathrm{d}\tau}\cdot\bm{\nabla}_{w}\right)\mathcal{P},\\ (\bm{\nabla}\bm{\nabla}\mathcal{P})^{(0)}&=\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\left(\hat{\bm{x}}\frac{\partial}{\partial w_{x}}+\hat{\bm{y}}\frac{\partial}{\partial w_{y}}\right)\mathcal{P}=\bm{\nabla}_{w}\bm{\nabla}_{w}\mathcal{P}.\end{split}$ (14)
We have ordered our terms such that the lowest order derivative of $\mathcal{P}$ with respect to the beam width is of the length scale $\sim W$, whilst the point-to-point variation in our profile must be of the typical length scale of the system $\sim L$. Explicitly:
$\frac{|(\bm{\nabla}\mathcal{P})^{(0)}|}{\mathcal{P}}\sim\frac{1}{W}\quad\mathrm{and}\quad\left|\frac{\partial\mathcal{P}}{\partial\tau}\right|\sim\frac{g}{L}{\mathcal{P}}.$ (15)